satellite73 (satellite73):

A somewhat easier probability problem. Adam and Eve take turns flipping coins. The first one to get heads wins. If Eve goes first, what is the probability that she wins?

OpenStudy (anonymous):

answer is 2/3

OpenStudy (anonymous):

because the probability eve will win is P(H)+P(TTH)+P(TTTTH)+P(TTTTTTH)+P(TT....H) this is a geometric sereis .. so u can solve it and answer is 2/3

OpenStudy (anonymous):

yes. can you find a solution without summation?

OpenStudy (anonymous):

yes... because the sum willbe a/(1-r) where a=1/2 and r=1/2^2

OpenStudy (anonymous):

i guess i meant a method that does not require geometric series

OpenStudy (anonymous):

I can solve it using this recurence, \[ T=\frac12+ \frac14T \] The chance of the Eve winning on her first turn is \( frac12 \), or she has a \( \frac14 \) chance of getting to her next turn, and like this follows , now...

OpenStudy (anonymous):

This much simpler than solving an geometric infinite series, isn't? ;-)

OpenStudy (anonymous):

and the reason i like this problem so much is that if you look closely, the method ffm wrote is exactly the method and reasoning one uses to derive the formula for the sum of a geometric series. the algebra is identical

OpenStudy (anonymous):

sat, do you have more of these in your sleeves?

OpenStudy (anonymous):

how about this one (similar reasoning) Adam and Eve roll a die. if it shows 1 or 2, Eve wins an apple from Adam and if it shows 3, 4, 5, 6 Adam wins and apple from Eve. Eve has two apples, Adam has only one. What is the probability Adam gets both Eve's apples before he loses his.

OpenStudy (anonymous):

answer is 3/7 ....

OpenStudy (anonymous):

check the algebra i got something similar but different

OpenStudy (anonymous):

wt did u get ? 6/7

OpenStudy (anonymous):

no, but it is early so i could be wrong. method?

OpenStudy (anonymous):

sat, teacher to student advice, .. I am thinking to spend some time reading this chapter : http://athenasc.com/Prob-2nd-Ch1.pdf precisely the problems is it worthy ?

OpenStudy (anonymous):

@ffm when you wrote "easier than solving geomtric series" the hook is easier but the algebra is identical. \[x=\frac{1}{2}+\frac{1}{4}x\] \[x-\frac{1}{4}x=\frac{1}{2}\] \[x(1-\frac{1}{4})=\frac{1}{2}\] \[x=\frac{\frac{1}{2}}{1-\frac{1}{4}}\]

OpenStudy (anonymous):

aha cute!

OpenStudy (anonymous):

but easier I mean you don't have to go through the steps of identifying the series -> remembering the formula ( or deriving it if you can't remember ) but I can see now that thte algebra is pretty much akin.

OpenStudy (anonymous):

look romeo and juliet!

OpenStudy (anonymous):

not only is the algebra the same, but the reasoning is exactly the same. think about how you derive formula for solving a geometric series.

OpenStudy (anonymous):

Yes I saw that hehe :D now please advice me what do you think about the over all chapter.

OpenStudy (anonymous):

it looks good if you have seen the topic before, and a nice condensation. if you have not seen the topic before i would start with something more expansive

OpenStudy (anonymous):

No, I have completed a undergrad level course on probability, but I am quite baffled to see that I can't solve quite a few problems on first reading.

OpenStudy (anonymous):

sat, I usually derive it for n terms and then take the limit to infinity

OpenStudy (anonymous):

then again you are mathematically mature so you can probably get a lot out of this.

OpenStudy (anonymous):

what i meant by reasoning is the same is this. usually you start with \[x =a+ ar + ar^2 +ar^3 + ...\] and then say "multiply by r and get \[rx =ar+ ar^2+ar^3+...\] so \[x-rx =a\] and therefore \[x=\frac{a}{1-r}\]and you see that this is precisely your reasoning when you wrote \[T=\frac{1}{2}+\frac{1}{4}T\] \[T-\frac{1}{4}T=\frac{1}{2}\] etc

OpenStudy (anonymous):

Cute, isn't this something that we can use to prove that 0.99999999999...= 1?

OpenStudy (anonymous):

this is the probabilistic "proof" for the algebra solution to a geometric series. of course in probabiltiy we don't have to worry about convergence because \[p<1\] (unless of course p = 1)

OpenStudy (anonymous):

I got it, it's the standard difference trick used in many places :)

OpenStudy (anonymous):

sure you can write \[\overline{.9}\] as \[\frac{9}{10}+\frac{9}{10^2}+...\]and sum it up \[\frac{\frac{9}{10}}{1-\frac{1}{10}}=1\]

OpenStudy (anonymous):

Great sat, I never thought about probabilistic proof! Thanks a lot :D

OpenStudy (anonymous):

yw btw i got \[\frac{4}{7}\] for the answer to second question maybe i will post it again

OpenStudy (anonymous):

New thread please :)